A student puts $10,000 in a savings account that pays 16% annual interest, compounded semi-annually and quarterly. Round to the nearest cents.

a) How much money will the student have at the end of 5 years? Ending Value: $

b) How much interest will the student have earned in 5 years? Interest Earned: $

c)What is the effective yield (APY)? APY: %, (written as percent, round to the two decimal place as needed)
Answer #1
Particulars a. Semi-annual b. Qtrly
Deposit              10,000              10,000
× FVF            2.15892            2.19112
Future value (answer)        21,589.25        21,911.23

c

APY

Semi-annual = (1+16%/2)^2 -1 = 16.64%

Quarterly = (1+16%/4)^4-1 = 16.99%